# Smooth Tests for Normality in ANOVA

### v2

@@ Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is rarely formally tested in practice. In this paper, we propose Neyman’s smooth tests for assessing the normality assumption across various types of ANOVA models. The proposed test statistics are constructed based on the Gaussian probability integral transformation of ANOVA residuals. Under the null hypothesis of normality, the test statistics are asymptotically Chi-square distributed, with degrees of freedom determined by the dimension of the smooth model (the number of orthonormal functions). A data-driven selection of the model dimension using a modified Schwarz’s criterion is also discussed. Simulation studies demonstrate the effectiveness of our proposed method.
@@ 1 Introduction
Analysis of Variance (ANOVA) is an important and widely used method in exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA models has been well established and developed in the literature, see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), Montgomery, (2017). A standard assumption in ANOVA models is that the random errors are normally distributed. Nonnormality may make the normal-based inference invalid. In terms of estimation, departures from normality generally undermine the validity of the normal-theory variance formulae for point estimates of variance components, because these formulae rely on theoretical results that represent the mean squares of random effects as constant multiples of Chi-square variables (Scheffé,, 1959). From the view of testing, the $F$-test employed in ANOVA is sensitive to nonnormality if the group sizes are unequal (Ali and Sharma,, 1996). The effects of nonnormality on size and power of the $F$-test are also extensively investigated in details (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964, 1971; Donaldson,, 1968; Janic-Wróblewska and Ledwina,, 2000). Therefore, it is crucial to evaluate the normality of ANOVA models.
@@
While a broad range of normality tests have been proposed for observed data and linear models, comparatively little effort has been devoted to systematically checking the normality assumption in ANOVA. A direct approach to normality diagnosis is to plot a normal probability plot of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which is intuitive but lacks a theoretical guarantee. To our knowledge, formal hypothesis testing procedures for assessing normality in ANOVA appear to be relatively scarce in the existing literature. Existing methods are merely a simple application of classical normality tests to ANOVA residuals, see, for instance, Bonett and Woodward, (1990), Bonett and Seier, (2002) and Hwang and Wei, (2006). Although applicable in practice, these methods do not precisely provide a rigorous assessment of normality within the ANOVA framework, nor do they account for the specific structure and constraints inherent in ANOVA models.
@@
In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing the normality assumption in various types of ANOVA models. Neyman’s smooth test has gained popularity across various scientific disciplines due to its well-known advantages in both theoretical and practical aspects. It has proven to be powerful in detecting against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022). Specifically, we first re-express the problem of testing for normality as testing for uniformity via the Gaussian probability integral transformation (PIT), and then construct the test statistics through the Gaussian PIT of ANOVA residuals. Under mild conditions for each model, the proposed test statistics are asymptotically Chi-square distributed under the null. Furthermore, we analyze the power properties of the tests under fixed and local alternatives. A modified Schwarz’s selection rule is proposed to determine the order of tests. Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
@@
The advantages and contributions of our proposed tests are as follows. First, the proposed asymptotic Chi-square tests are simple to implement, in the sense that the critical values can be obtained directly from the limiting Chi-square distributions without bootstrapping. Second, we systematically investigate the properties of test statistics for three types of one-way fixed effects models with potentially diverging numbers of groups in a unified asymptotic framework. The parameter estimation effects in each model are well interpreted through our analysis procedure. The heterogeneity in group means and group variances does contribute to the specific form of test statistics and to the corresponding regularity conditions, but does not alter the convergence rate, which is only related to the total sample sizes under both the null and the alternatives. Lastly, we clarify the motivation and theoretical properties of data-driven smooth tests under ANOVA settings, and provide a revised limiting null distribution of the data-driven test statistics in finite samples.
The rest of this paper is organized as follows. Section 2 presents the general testing framework. Section 3 establishes the theoretical results for smooth tests in three types of ANOVA models. In Section 4, a data-driven approach is proposed to determine the order of tests. Simulation results in Section 5 demonstrate good finite-sample performance of our proposed methodology. Some extensions are discussed in Section 6. Section 7 concludes. The proof of theoretical results and some practical issues are gathered in the supplementary material.
@@ 2 The testing framework
Our problem of interest is whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$ (denoted as $\mathcal{N}(0,\sigma^{2})$) for some $\sigma>0$. We aim to test for the null hypothesis $H_{0}:e\sim\mathcal{N}(0,1)$ against the alternatives of nonnormality. To motivate our methodology, consider the transformed random variable $Z=\Phi(e)$, with $\Phi(\cdot)$ denoting the CDF of the standard normal distribution. Under the null hypothesis of normality, the CDF of $Z$ is given by
@@
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1$ for $z\in[0,1]$. That is, $Z\sim\mathcal{U}[0,1]$ under $H_{0}$, where $\mathcal{U}[0,1]$ marks the uniform distribution on the interval $[0,1]$. Under the alternatives, the PDF $g(z)$ differs from 1. The distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the basis for Neyman’s smooth test. Neyman, (1937) considered the following smooth alternative to the uniform density:
@@
where $c(\bm{\theta}_{K})$ is a constant of integration depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ is an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
@@
The null hypothesis that $Z\sim\mathcal{U}[0,1]$ can then be tested by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
If we had independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
@@
where $K$ is some fixed positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$. Under $H_{0}$, $n\Psi_{K}^{2}$ converges asymptotically to a Chi-square distribution with $K$ degrees of freedom (denoted as $\chi_{K}^{2}$), and the test has the local optimal properties of Rao’s score test.
However, it is always the case that $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We may use some approximation $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and some estimator $\hat{\sigma}$ for $\sigma$ (e.g., the residuals and the sample standard deviation, respectively). Then, consider the following approximation for $Z_{i}$:
$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n.$
With $\{\hat{Z}_{i}\}_{i=1}^{n}$ replacing $\{Z_{i}\}_{i=1}^{n}$, we can construct a feasible smooth test statistic of the following form:
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$
Due to the replacement of $Z_{i}$ by $\hat{Z}_{i}$, the sequence $\{\hat{Z}_{i}\}_{i=1}^{n}$ may not be i.i.d. any longer. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), $\hat{Z}_{i}$ has non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the above form of the test statistic and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates us to study the estimation effects carefully and then derive the correct quadratic-form test statistics.
It is also important to address the issue of sample sizes in ANOVA models. Throughout this section, to highlight the essence of smooth tests, we do not directly consider the ANOVA models but first examine the testing for a single random sequence of sample size $n$. This simplified setting captures the core ideas of the methodology and facilitates the derivation of key asymptotic properties. In the complete ANOVA framework, additional complications arise due to multiple sample sizes and the indexing structure, which require a more elaborate treatment. Besides, from a practical perspective, our method is expected to accommodate scenarios in which the number of groups grows with the total sample size. In subsequent analysis, we will systematically tackle these issues for specific ANOVA models.
@@ 3 Testing for normality in one-way fixed effects models
In this section, we establish smooth tests for normality in three types of one-way fixed effects models where either the group means or the group variances are equal. We examine the testing procedure and the theoretical properties separately for each case. Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$. Define $N=\sum_{j=1}^{J}N_{j}$ as the total sample size. In our theoretical framework, the number of groups $J$ is allowed to diverge as $N$ increases.
@@ 3.1 Same mean and same variance
We first investigate the ANOVA models with equal group means and equal group variances. Consider the following model:
@@
where $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. random errors with mean zero and variance one, $\mu$ is the common group mean, and $\sigma$ is the standard deviation of random errors. We also define $Z_{ij}=\Phi(e_{ij})$.
We estimate $\mu$ and $\sigma^{2}$ in (3.1) by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To derive the test statistic and its asymptotic distribution, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. The following two assumptions are imposed for our asymptotic analysis.
@@
$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, variance one, and finite fourth moments.
@@
For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The two mild conditions are in line with similar assumptions adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). Introduce two constants:
@@
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is asymptotically equivalent to the sum of three terms: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
With the assistance of Theorem 1, by the central limit theorem (CLT), we can obtain the asymptotic normality of the $K$-dimensional vector
@@
where $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
@@
Based on this fact, our proposed feasible smooth test statistic is defined as
@@
Corollary 1 demonstrates an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
@@
Now we turn to the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives and a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
@@
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed, which leads to different asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ from those under $H_{0}$. To formally present the corresponding theoretical results under $H_{1}$, we need some notations. Let
@@
for $k=1,\ldots,K$. Note that under $H_{0}$, $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ due to $f(x)=\phi(x)$. The following theorem demonstrates the asymptotic properties under $H_{1}$.
@@
From Theorem 2, one can see that under $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
@@
the smooth test statistic $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, and the asymptotic power of the test is 1. Specifically, the power function is given by
@@
For further power analysis, we consider a Pitman-type sequence of local alternatives converging to the null hypothesis at a suitable rate, which is specified as follows:
@@
where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Denote
@@
The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the local power of the test.
@@
It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to have a nontrivial power against the Pitman sequence of local alternatives. It would be impossible to distinguish the alternatives from the null if $\delta_{N}=o(N^{-1/2})$. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ has nontrivial asymptotic power against the local alternatives.
@@
We should emphasize that the proposed smooth test is not a “consistent” test in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power go to 1. Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, if $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ such that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and we take the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$, then for $Z=\Phi(e)$, we have
@@
which leads to $\mathrm{E}[\pi_{1}(Z)]=0$. In fact, smooth tests are neither directional nor omnibus; that is, they maintain decent power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for details.

### v3

@@ Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is seldom subjected to formal testing in practice. In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models. The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals and are shown to follow an asymptotic Chi-square distribution under the null hypothesis, with degrees of freedom determined by the dimension of the smooth model. We further propose a data-driven selection of the model dimension based on a modified Schwarz’s criterion. Monte Carlo simulations demonstrate that the tests maintain the nominal size and achieve high power against a wide range of alternatives. Our framework thus provides a systematic and effective tool for formally validating the normality assumption in ANOVA models.
@@ 1 Introduction
Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA has been well established in the literature; see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), and Montgomery, (2017). A standard assumption in ANOVA is that the random errors are normally distributed. Violations of this assumption may invalidate normal-based inference. From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (Scheffé,, 1959). From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sizes are unequal (Ali and Sharma,, 1996). The impact of nonnormality on the size and power of the $F$-test has been extensively studied (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964; Donaldson,, 1968; Tiku,, 1971; Janic-Wróblewska and Ledwina,, 2000). These considerations highlight the importance of rigorously evaluating the normality assumption in ANOVA models.
@@
Although a variety of normality tests have been proposed for observed data and linear models, relatively little attention has been paid to systematically assessing normality in ANOVA. A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which are intuitive but lack theoretical justification. To our knowledge, formal hypothesis testing procedures designed explicitly for ANOVA are scarce in the literature. Most existing methods simply apply classical normality tests to ANOVA residuals, see, for example, Bonett and Woodward, (1990), Bonett and Seier, (2002), and Hwang and Wei, (2006). While applicable in practice, such approaches do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the structural constraints inherent in ANOVA models.
@@
In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing normality in various types of ANOVA models. Neyman’s smooth test is widely used in diverse scientific fields due to its theoretical soundness and practical effectiveness, with proven power against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022). Specifically, we reformulate the normality testing problem as one of uniformity via the Gaussian probability integral transform (PIT), and construct test statistics based on the Gaussian PIT of ANOVA residuals. Under mild conditions, the proposed statistics are asymptotically Chi-square distributed under the null hypothesis. We further analyze the power properties under both fixed and local alternatives. A modified Schwarz’s selection rule is proposed to determine the direction of tests. Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
@@
The main contributions of this paper are threefold. First, the proposed asymptotic Chi-square tests are simple to implement: critical values are obtained directly from the limiting distributions without resorting to resampling. Second, we systematically investigate the properties of the test statistics in three types of one-way fixed effects models, potentially with a diverging number of groups, within a unified asymptotic framework. Our analysis clarifies how parameter estimation affects the test statistics: heterogeneity in group means and variances alters their explicit form and the corresponding regularity conditions, but the convergence rate depends only on the total sample size under both null and alternative hypotheses. Third, we provide theoretical justification for data-driven smooth tests in the ANOVA setting, including a revised limiting null distribution for the corresponding statistics in finite samples.
The remainder of this paper is organized as follows. Section 2 introduces the general testing framework. Section 3 establishes theoretical results for three types of ANOVA models. Section 4 proposes a data-driven testing procedure. Section 5 reports simulation results. Section 6 discusses further extensions. Section 7 concludes. Proofs and additional numerical results are provided in Appendix A.
@@ 2 The testing framework
Our problem of interest is to determine whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$. Equivalently, this can be formulated as testing the null hypothesis about the standardized random error $H_{0}:e\sim\mathcal{N}(0,1)$ against alternatives of nonnormality. To motivate our methodology, consider the transformation $Z=\Phi(e)$, where $\Phi(\cdot)$ denotes the cumulative distribution function (CDF) of the standard normal distribution. Under the null hypothesis of normality, the CDF of the transformed random variable $Z$ is given by
@@
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1,z\in[0,1]$, that is, $Z\sim\mathcal{U}[0,1]$, where $\mathcal{U}[0,1]$ denotes the uniform distribution on the interval $[0,1]$. Under the alternative, the density $g(z)$ deviates from unity. This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the foundation for Neyman’s smooth test. In particular, Neyman, (1937) introduced the following smooth alternative to the uniform density:
@@
where $c(\bm{\theta}_{K})$ is a a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
@@
The null hypothesis $Z\sim\mathcal{U}[0,1]$ can thus be assessed by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
@@
where $K$ is a fixed and given positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$. Under $H_{0}$, the statistic $n\Psi_{K}^{2}$ converges in distribution to a Chi-square law with $K$ degrees of freedom, denoted by $\chi_{K}^{2}$. Moreover, the test inherits the local optimal properties of Rao’s score test.
However, in practice, the sequence $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We instead rely on approximations $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\hat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively). Accordingly, we consider the following approximation for $Z_{i}$:
$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n,$
and construct a feasible test statistic of the following form:
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$ (2.4)
The random variables $\{\hat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d. any longer due to estimation effects caused by $\hat{\varepsilon}_{i}$ and $\hat{\sigma}$. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\hat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the form of the “naive” statistic (2.4) and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates a detailed analysis of estimation effects, leading to the correct quadratic-form test statistic.
Another important issue concerns sample sizes in ANOVA models. To highlight the essence of smooth tests, we first focus on the simplified setting of a single random sequence with sample size $n$ throughout this section. This simplified setting captures the core methodological ideas and facilitates the derivation of key asymptotic properties. In contrast, the complete ANOVA setting involves additional complications arising from multiple sample sizes and the associated indexing structure, which demand more refined analysis. From a practical standpoint, our method is further designed to accommodate scenarios where the number of groups increases with the total sample size. These issues will be systematically investigated in subsequent analysis of specific ANOVA models.
@@ 3 Testing for normality in one-way fixed effects models
In this section, we develop smooth tests for normality in three types of one-way fixed effects models, where each model assumes equality in at least one of the group means or the group variances. The testing procedure and its theoretical properties are investigated separately for each case. Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$. The total sample size is defined as $N=\sum_{j=1}^{J}N_{j}$. In our theoretical framework, the number of groups, $J$, is allowed to diverge as $N$ increases.
@@ 3.1 Same mean and same variance
We begin with the ANOVA models where both the group means and group variances are equal. Specifically, consider the model
@@
where $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. standardized random errors with mean zero and unit variance, $\mu$ denotes the common group mean, and $\sigma$ is the standard deviation of random errors $\{\varepsilon_{ij}\}_{i=1,j=1}^{N_{j},J}$. We also define $Z_{ij}=\Phi(e_{ij})$.
The parameters $\mu$ and $\sigma^{2}$ in (3.1) are estimated by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ for $k=1,\ldots,K$. For this purpose, we impose the following two assumptions.
@@
$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, unit variance, and finite sixth moments.
@@
For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice continuously differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). For subsequent analysis, we introduce two constants:
@@
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order term: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
With the assistance of Theorem 1, and by invoking the central limit theorem (CLT), we obtain the asymptotic normality of the $K$-dimensional vector
@@
where the asymptotic covariance matrix $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
@@
Based on this result, we define the feasible smooth test statistic as
@@
Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
@@
We now investigate the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
@@
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed. Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ differs from that under $H_{0}$. To formally state the corresponding theoretical results under $H_{1}$, we introduce the following notations. Let
@@
for $k=1,\ldots,K$. Note that under $H_{0}$, the constants $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ since $f(x)=\phi(x)$. The following theorem characterizes the asymptotic properties of $\hat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.
@@
From Theorem 2, it follows that under the alternative hypothesis $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
@@
then the smooth test statistic satisfies $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$. Specifically, the power function is given by
@@
For further investigation of the test’s power, we consider a Pitman-type sequence of local alternatives that converges to the null hypothesis at an appropriate rate, which is specified as follows:
@@
where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Define
@@
The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the nontrivial local power of the test.
@@
It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to achieve nontrivial power against the Pitman-type sequence of local alternatives. If $\delta_{N}=o(N^{-1/2})$, it would be impossible to distinguish the alternatives from the null. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
@@
We emphasize that the aforementioned smooth test is not “consistent” in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$. Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$. For $Z=\Phi(e)$, we have
@@
which implies $\mathrm{E}[\pi_{1}(Z)]=0$. In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for further discussion.